Part I. Prepare and Test Standard Solutions 1. Obtain and wear goggles. 2. Label four small beakers 1-4. Obtain small volumes of 0.200 M Fe(NO3)3, 0.0020 M SCN-, and distilled water. Prepare four solutions according to the chart below Use graduated cylinders to measure the solutions. Mix each solution thoroughly Measure and record the temperature of either of the solutions - remember that the equilibrium constant (Kea) depends on temperature. Dont cross-contaminate the solutions. Technical note 1: The Fe(NO3)3 solutions in this experiment are prepared in 1.0 M HNO3 (to prevent the decomposition of Fe3*) and should be handled with care Technical note 2: The mixtures you will prepare are light sensitive. You need to measure the absorbances of these four mixtures within 2-5 minutes of preparing them. Beaker 0.0020 M KSCN 0.200 M Fe(NO3)3 (mL) 5.0 5.0 5.0 5.0 (mL) 41.0 42.0 43.0 44.0 4.0 3.0 2.0 2 4 3. Prepare a blank by filling a cuvette 3/4 full with 0.200 M Fe(NO3)3 solution. To correctly use cuvettes, remember: . Wipe the outside of each cuvette with a lint-free tissue. Handle cuvettes only by the top edge of the ribbed sides. . Dislodge any bubbles by gently tapping the cuvette on a hard surface. . Always position the cuvette so that the light passes through the clear sides. Measure absorbance of the blank cuvette with a spectrophotometer station (see back of the manual for tips on spectrophotometers). Then measure all the four samples, one by one, at the same wavelength. Write down the absorbance values. 4.
Answer #1

There's various ways to this, but let me tell you the simpliest one, first the overall reaction would be:

Fe(NO3)3(ac) + KSCN(ac) -------> KNO3(ac) + (FeSCN)(NO3)2(ac)

We know then that we have a 1:1 relation on the reactants and products. Let's calculate moles for each reactant:

moles Fe(NO3)3:

1. moles = 0.2 * 0.005 = 1x10-3 moles

moles of KSCN:

1. moles = 0.002 * 0.004 = 8x10-6 moles

2. moles = 0.002 * 0.003 = 6x10-6 moles

3. moles = 0.002 * 0.002 = 4x10-6 moles

4. moles = 0.002 * 0.003 = 2x10-6 moles

This means that the limitant reactant is the KSCN so, the moles obtained for Fe(SCN)2+ would be the number of moles of SCN, so, the concentration would have to be:

1. [Fe(SCN)2+] = 8x10-6 moles / 0.050 = 1.6x10-4 M

2. [Fe(SCN)2+] = 6x10-6 moles / 0.050 = 1.2x10-4 M

3. [Fe(SCN)2+] = 4x10-6 moles / 0.050 = 8.0x10-5 M

4. [Fe(SCN)2+] = 2x10-6 moles / 0.050 = 4.0x10-5 M

For the part of the graph, all you have to do is plot in an excel sheet (or in your calculator) Absorbance vs Concentration (absorbance in y axis and concentration in x axis) Doing this, the linear equation obtained would be:

y = 5605x - 0.1465

Now, are you sure that the 3 values of absorbance of 3 and 4 are those? because when I entered those values, the r2 = 0.4561 and this is not linear. The absorbance would have to higher in 3 and 4 not lower. Could you please check that?. When you do that, just replace the values and get the new linear equation.

Tell me in a comment if there's something else you need or something to be fixed.

Hope this helps